OpenStudy (anonymous):

Can some please help me put this in standard form? I'm pretty sure it's a parabola, but I'm confused on how to get this tin standard. y^2+4x-6y+5=0

OpenStudy (horsegirl27):

well standard form is A+B=C

OpenStudy (daniellelovee):

I would say the standard form is 3+2sqrt-x+1

OpenStudy (daniellelovee):

=y

OpenStudy (daniellelovee):

\[y=3\pm2\sqrt{-x+1}\]

OpenStudy (horsegirl27):

Ax+By=C and C will be 5

OpenStudy (anonymous):

you guys mind helping me? 1/2 + 3/2d = -1/d I got 11

OpenStudy (horsegirl27):

4x-5y^2=5

OpenStudy (horsegirl27):

@Avonniculley post it as a different question